# Asymptotic structure. III. Excluding a fat tree

### v1

@@ 1 Introduction
(In [10], we proved a version of this with path-width instead of line-width, and 2k12k-1 replaced by $k$, but Hickingbotham tells us that his different method of proof yields the result 1.5, and he will add it to his paper.) Thus, to complete the proof of 1.4, it suffices to prove the following:
@@ 4 Realms
  • For each fort $X$ of $\mathcal{T}$, $\operatorname{bd}(X)$ has quasi-bound at most $(\alpha,\beta)$;
@@ 6 Passages
Let $G$ be a graph that does not contain a $c$-superfat $H_{\ell}$-minor, and let $\mathcal{T}$ be an optimal $k$th-century realm in $G$, and let $S=\bigcup_{X\in\mathcal{T}^{0}}\operatorname{\Delta}_{\mathcal{T}}(X)$. Then $S$ has quasi-line-width at most $(5\alpha,\beta+d_{0})$.
Proof.  Let $J$ be the graph with vertex set the set of forts of $\mathcal{T}$ together with the set of all maximal communities of $\mathcal{T}$, where we say $X,Y\in V(J)$ are adjacent if $X$ adjoins $Y$ (and consequently at least one of $X,Y$ is a fort). From 6.2, each maximal community in $J$ has degree at most two in $J$, and its neighbours in $J$ are forts. By 6.2, $J$ has line-width at most three. For each $X\in V(J)$, if $X$ is a fort of $\mathcal{T}$ let $A(X)=\operatorname{\Delta}_{\mathcal{T}}(X)$, and if $X$ is a maximal community of $\mathcal{T}$, let $A(X)$ be the union of the sets $\operatorname{\Delta}_{\mathcal{T}}(Y)\;(Y\in X)$. Thus the sets $A(X)\;(X\in V(J))$ are pairwise disjoint and have union $S$. For each $X\in V(J)$, $A(X)$ has quasi-line-width at most $(\alpha,\beta+d_{0})$, and its boundary has quasi-size at most $(\alpha,\beta+d_{0})$. By 2.2, $S$ has quasi-line-width at most $(5\alpha,\beta+d_{0})$. This proves 6.3.
@@ 8 Revolution
Since $W$ has quasi-line-width at most $(5\alpha,\beta+d_{0})$ by 6.3, and each member of $B$ has quasi-line-width at most $(8\alpha,\beta+(\ell-k-1)d_{0})$, and there are no edges between any two of the sets in $B\cup\{W\}$, it follows that $W\cup V(B)$ has quasi-line-width at most
$\left(8\alpha,\beta+(\ell-k-1)d_{0}\right).$
@@
$\left(8\cdot 3^{\ell-k-2}\alpha+56(\ell-k-1)3^{\ell-k-2}\alpha,\beta+2(\ell-k-1)d_{0}\right)$
@@
$\left(7(\ell-k)3^{\ell-k}\alpha,\beta+3(\ell-k-1)d_{0}\right).$
@@
Let $\mathcal{T}$ be an optimal $k$th-century realm, and let $\mathcal{G}$ be a stable small government for $\mathcal{T}$. Let $C$ be a set of rebels, in communication and maximal with this property; and let $W$ be the union of the sets $L_{\mathcal{G}}(X)\;(X\in C)$. Then $W$ has quasi-line-width at most $(5\alpha,\beta+d_{0})$, and $\operatorname{bd}(W)$ has quasi-size at most
@@ 9 A new century
    For each fort $X$ of $\mathcal{T}^{\prime}$, $\operatorname{bd}(X)$ has quasi-bound at most $(\alpha_{k+1},\beta_{k+1})$.

### v2

@@ 1 Introduction
Thus, to complete the proof of 1.4, it suffices to prove the following:
@@ 4 Realms
  • Each fort of $\mathcal{T}$ has quasi-bound at most $(\alpha,\beta)$;
@@ 6 Passages
Let $G$ be a graph that does not contain a $c$-superfat $H_{\ell}$-minor, and let $\mathcal{T}$ be an optimal $k$th-century realm in $G$, and let $S=\bigcup_{X\in\mathcal{T}^{0}}\operatorname{\Delta}_{\mathcal{T}}(X)$. Then $S$ has quasi-line-width at most $(10\alpha,\beta+d_{0})$.
Proof.  Let $J$ be the graph with vertex set the set of forts of $\mathcal{T}$ together with the set of all maximal communities of $\mathcal{T}$, where we say $X,Y\in V(J)$ are adjacent if $X$ adjoins $Y$ (and consequently at least one of $X,Y$ is a fort). From 6.2, each maximal community in $J$ has degree at most two in $J$, and its neighbours in $J$ are forts. By 6.2, $J$ has line-width at most three. For each $X\in V(J)$, if $X$ is a fort of $\mathcal{T}$ let $A(X)=\operatorname{\Delta}_{\mathcal{T}}(X)$, and if $X$ is a maximal community of $\mathcal{T}$, let $A(X)$ be the union of the sets $\operatorname{\Delta}_{\mathcal{T}}(Y)\;(Y\in X)$. Thus the sets $A(X)\;(X\in V(J))$ are pairwise disjoint and have union $S$. For each $\in V(J), XX$ has quasi-bound at most $(a,b)$. Every vertex in $\operatorname{\Delta}_{\mathcal{T}}(X)$ either belongs to $X$ or has distance at most $d_{0}$ from $\operatorname{bd}(X)$, and in particular, every vertex in $\operatorname{bd}(\operatorname{\Delta}_{\mathcal{T}}(X))$ has distance at most $d_{0}$ from $\operatorname{bd}(X)$. Since $\operatorname{bd}(X)$ has quasi-size at most $(a,b)$, it follows that $\operatorname{bd}(\operatorname{\Delta}_{\mathcal{T}}(X))$ has quasi-size at most $(a,b+d_{0})$. Similarly, $\operatorname{\Delta}_{\mathcal{T}}(X)\setminus X$ has quasi-size at most $(a,b+d_{0})$, and so, by adding the set $\operatorname{\Delta}_{\mathcal{T}}(X)\setminus X$ to each bag of a line-decomposition of $X$, we deduce that $\operatorname{\Delta}_{\mathcal{T}}(X)$ has quasi-line-width at most $(2a,b+d_{0})$. By 2.2, $S$ has quasi-line-width at most $(10\alpha,\beta+d_{0})$. This proves 6.3.
@@ 8 Revolution
Since $W$ has quasi-line-width at most $(10\alpha,\beta+d_{0})$ by 6.3, and each member of $B$ has quasi-line-width at most $(8\alpha,\beta+(\ell-k-1)d_{0})$, and there are no edges between any two of the sets in $B\cup\{W\}$, it follows that $W\cup V(B)$ has quasi-line-width at most
$\left(10\alpha,\beta+(\ell-k-1)d_{0}\right).$
@@
$\left(10\cdot 3^{\ell-k-2}\alpha+56(\ell-k-1)3^{\ell-k-2}\alpha,\beta+2(\ell-k-1)d_{0}\right)$
@@
$\left(7(\ell-k)3^{\ell-k}\alpha,\beta+2(\ell-k-1)d_{0}\right).$
@@
Let $\mathcal{T}$ be an optimal $k$th-century realm, and let $\mathcal{G}$ be a stable small government for $\mathcal{T}$. Let $C$ be a set of rebels, in communication and maximal with this property; and let $W$ be the union of the sets $L_{\mathcal{G}}(X)\;(X\in C)$. Then $W$ has quasi-line-width at most $(10\alpha,\beta+d_{0})$, and $\operatorname{bd}(W)$ has quasi-size at most
@@ 9 A new century
    Each fort of $\mathcal{T}^{\prime}$ has quasi-bound at most $(\alpha_{k+1},\beta_{k+1})$.